# you are in the wrong room

### v1

@@ Abstract.
If taken seriously, the advice in the title leads to interesting combinatorics. Consider $N$ people moving between $M$ rooms as follows: at each step, simultaneously, the smartest person in each room moves to a different room of their choice. There are $M^{N}$ possible configurations. We classify the pairs $(N,M)$ such that every configuration can be reached from every other. More precisely, let $G(N,M)$ be the directed graph with vertices representing configurations and edges representing legal moves. We prove that the graph $G(N,M)$ is strongly connected if and only if $M\geq N+1$ (one extra room for maneuvering is both required and sufficient). In the case $M\leq N$, we explicitly describe the components. Proofs are based on algebraic properties of derangements.
@@ 1. Introduction
It is natural to ask whether any configuration can be reached from any other in a finite number of steps. The following theorem shows that having one extra room for maneuvering is both necessary and sufficient.
@@
By Theorem 1.1, for $M\leq N$ the graph $G(N,M)$ is not strongly connected. The next theorem shows that, in such case, a positive fraction of vertices forms a single strongly connected component (a giant component), while the remaining vertices have zero in-degree and positive out-degree. In other words, every directed path enters the giant component in at most one step and never leaves it afterwords.
@@
Figure 1. The configuration graph $G(4,3)$ for $N=4$ people and $M=3$ rooms. The set $V_{s}$ is labeled $\bullet$, the set $V_{c}$ is labeled $\blacksquare$, the other vertices are labeled $\times$. The vertices labeled $\bullet$ and $\times$ form a single strongly connected component. Every vertex labeled $\blacksquare$ has zero in-degree and form a strongly connected component on its own.
@@
We conclude the introduction with an open problem. Suppose each room’s smartest person independently choses the next room uniformly at random among the other rooms. Then the next vertex in $G(N,M)$ is chosen uniformly at random among the out-neighbors. While Theorem 1.1 and Theorem 1.2 classify recurrent vertices in $G(N,M)$, they do no determine their asymptotic frequencies. This is left as an open problem. For a survey on random walks on graphs, see [4]. Note that the random walk on $G(N,M)$ can also be viewed as $N$ random walkers moving among $M$ rooms under a priority rule. While multiple random walks have been extensively studied [1], such a priority rule seems to be novel.
@@
The statement is vacuous for $n=1$ and trivial for $n=2$. Let $n\geq 4$. Husemoller [3, Proposition 4] attributes to Gleason the result that every even permutation is product of two $n$-cycles (see [5] for a more general result). In particular, every even permutation is product of $2$ derangements. Since every odd permutation is product of an even permutation and a transposition, it remains to show that every transposition is product of $2$ derangements.

### v2

@@ Abstract.
If taken seriously, the advice in the title leads to interesting combinatorics. Consider $N$ people moving between $M$ rooms as follows: at each step, simultaneously, the smartest person in each room moves to a different room of their choice, while no one else moves. The process repeats. Let $G(N,M)$ be the directed graph with vertices representing the $M^{N}$ possible configurations and edges representing possible moves. We classify the pairs $(N,M)$ such that every configuration can be reached from every other. More precisely, we prove that the graph $G(N,M)$ is strongly connected if and only if $M\geq N+1$ (one extra room for maneuvering is both required and sufficient). In the case $M\leq N$, we explicitly describe strongly and weakly connected components. Proofs are either impossibility arguments, or explicit algorithms based on subgroups generated by derangements.
@@ 1. Introduction
It is natural to ask whether any configuration can be reached from any other in a finite number of steps. In Section 2 we will prove the following theorem, showing that having one extra room for maneuvering is both necessary and sufficient.
@@
By Theorem 1.1, for $M\leq N$ the graph $G(N,M)$ is not strongly connected. The next theorem, proved in Section 3, shows that in such case a positive fraction of vertices forms a single strongly connected component (a giant component), while the remaining vertices have zero in-degree and positive out-degree. In other words, every directed path enters the giant component in at most one step and never leaves it afterwords.
@@
Figure 1. The configuration graph $G(4,3)$ for $N=4$ people and $M=3$ rooms. The set $V_{s}$ is labeled $\bullet$, the set $V_{c}$ is labeled $\blacksquare$, the other vertices are labeled $\times$. The vertices labeled $\bullet$ and $\times$ form a single strongly connected component. Every vertex labeled $\blacksquare$ cannot be reached from other vertices and forms a strongly connected component on its own.
@@
1.1. Towers of Hanoi
The problem discussed in this paper resembles the classical Towers of Hanoi problem [6, 2, 8]. There is a natural analogy between the two, with people corresponding to disks and rooms corresponding to towers. However, there are a few substantial differences. First, in the Towers of Hanoi problem, only one disk is moved at a time, while in the rooms problem, the number of people moving simultaneously equals the current number of non-empty rooms. Second, in the Towers of Hanoi problem, only the top disk of a tower can be moved (this is the smallest disk and the most recently placed in that tower), while in the room problem, such chronological order does not need to be preserved.
1.2. Random Walks
We conclude the introduction with an open problem. Suppose each room’s smartest person independently choses the next room uniformly at random among the other rooms. Then the next vertex in $G(N,M)$ is chosen uniformly at random among the out-neighbors. While Theorem 1.1 and Theorem 1.2 classify recurrent vertices in $G(N,M)$, they do no determine their asymptotic frequencies. This is left as an open problem. For a survey on random walks on graphs, see [5]. Note that the random walk on $G(N,M)$ can also be viewed as $N$ random walkers moving among $M$ rooms under a priority rule. While multiple random walkers have been extensively studied [1], such a priority rule seems to be novel.
@@ 1.2. Random Walks
The statement is vacuous for $n=1$ and trivial for $n=2$. Let $n\geq 4$. Husemoller [4, Proposition 4] attributes to Gleason the result that every even permutation is product of two $n$-cycles (see [7] for a more general result). In particular, every even permutation is product of $2$ derangements. Since every odd permutation is product of an even permutation and a transposition, it remains to show that every transposition is product of $2$ derangements.
@@
4. Acknowledgements
I would like to thank my colleague and friend Eric S. Vidal for pointing out the similarity to the Towers of Hanoi.